# TempMe: Video Temporal Token Merging for Efficient Text-Video Retrieval

### v1

@@ 3.2 Temporal Token Merging
To verify the efficiency of our proposed TempMe, we perform a detailed complexity analysis in Table 1. Following common protocols [36, 53], the frame length is set to 12 for MSR-VTT [49] and LSMDC [43], and to 64 for ActivityNet [27] and DiDeMo [1]. The complexity comparison between ToMe and our TempMe for these frame lengths is shown in Table 1. Compared to ToMe, our TempMe reduces tokens by more than 30% for both ViT-B/32 and ViT-B/16, effectively decreasing model complexity while significantly surpassing accuracy. Particularly, with a 12-frame length and ViT-B/16 backbone on MSRVTT, TempMe outputs only 5% of the input tokens, reaches 57% GFLOPs, and achieves a 5.3% R-Sum gain.
@@ 4 Experiments
Table 2: Comparisons in the text-to-video task on MSRVTT.
@@
Full Fine-tuning CLIP4Clip [36] 123.54 53.0 43.1 70.4 80.8 194.3 16.2
Parameter-Efficient Prompt [25] 0.08 58.2 40.4 66.3 77.3 184.0 16.7
Adapter [16] 0.26 53.1 41.9 69.9 78.7 190.2 14.9
LoRA [17] 0.49 53.0 43.7 68.9 80.4 193.0 16.0
PLEVU [23] 6.35 - 36.7 64.6 76.8 178.1 -
VoP [18] 14.10 58.0 44.7 70.5 79.2 194.4 16.2
DGL [53] 0.83 67.4 44.6 69.9 80.3 194.8 16.3
Parameter-Efficient & Inference-Efficient EVIT [32] 0.49 37.2 41.4 69.0 78.1 188.5 16.8
DiffRate [6] 0.49 36.8 41.5 68.6 79.8 189.9 16.3
STA [9] 0.49 35.7 42.6 69.5 78.8 190.9 17.0
ToMe [4] 0.49 40.2 42.9 68.3 80.2 191.4 16.2
TESTA [42] 0.59 40.6 43.7 69.0 79.4 192.1 16.8
@@
Parameter-Efficient MV-Adapter [22] 3.6 >210 46.2 73.2 82.7 202.1 -
RAP [5] 1.06 >210 46.5 73.9 82.0 202.4 12.1
VoP [18] 14.10 246.2 47.7 72.4 82.2 202.3 12.0
DGL [53] 0.83 251.2 48.3 71.8 80.6 200.7 13.4
@@ 4.1 Experimental Settings
Following common practice, we perform experiments on four widely used benchmarks for text-video retrieval: MSRVTT [49], ActivityNet [27], DiDeMo [1], and LSMDC [43]. Detailed descriptions of these datasets are provided in Appendix B.
@@
Following previous works [36, 23, 18, 53], we use the pre-trained CLIP as the backbone. We employ the AdamW optimizer [35] with a batch size of 128. For short video retrieval datasets like MSRVTT and LSMDC, the max word and frame lengths are set to 32 and 12. For long video retrieval datasets like ActivityNet and DiDeMo, the max word and frame lengths are set to 64 and 64. Unless noted otherwise, ViT-B/32 is employed as the backbone. More details are provided in Appendix B.
@@
We compare our TempMe with several strong baselines: (1) CLIP4Clip, a classic full fine-tuning method. For a fair comparison, we only consider the parameter-free type that does not introduce additional modules. (2) Prompt [25], LoRA [17], and Adapter [16], which are mainstream parameter-efficient fine-tuning methods. (3) PLEVU [23], VoP [18], DGL [53], MV-Adapter [22], and RAP [5], which are methods tailored for parameter-efficient text-video retrieval. (4) EVIT [32], DiffRate [6], and ToMe [4], which are token compression methods for image models. (5) STA [9] and TESTA [42], which are video token compression methods. For fairness, token compression methods apply LoRA like our TempMe (see Appendix A).
@@
Prompt 29.7 216.8 2369 194.2
Adapter 29.4 211.7 2364 196.8
LoRA 30.6 211.3 2364 201.4
VoP 25.0 246.2 2368 202.3
DGL 3.3 251.2 2416 200.7
DiffRate 40.6 138.5 588 202.4
ToMe 40.8 144.4 924 200.6
TempMe 45.1 121.4 127 206.7
@@
CLIP4Clip 211.3 1.00$\times$ 70.1GB 202.3
+ToMe 144.4 1.35$\times$ 61.6GB 201.6
+TempMe 121.4 1.57$\times$ 52.7GB 210.2
@@ 4.2 Comparisons with State-of-the-art Methods
In Table 2, we compare the t2v performance on MSRVTT. Our TempMe achieves significant improvements over previous methods, with a $\textbf{3.8}\%$ R-Sum increase in t2v using ViT-B/32 and a $\textbf{4.3}\%$ R-Sum increase in t2v using ViT-B/16, while maintaining minimal GFLOPs. Table 3 shows the t2v results on ActivityNet, DiDeMo, and LSMDC, demonstrating that TempMe consistently outperforms state-of-the-art methods. Detailed v2t results are provided in Appendix C.
@@ 4.3 Generalization Analysis
Table 6: Application in video foundation models for text-video retrieval. We conduct a evaluation of full fine-tuning on MSRVTT for the text-to-video task with UMT-B/16-25M.
@@
Table 7: Application in video foundation models for text-video QA. We conduct a evaluation of full fine-tuning on MSR-QA with UMT-B/16-25M.
@@
We extend our method to video foundation methods, such as UMT [31]. TempMe is built upon the text-image CLIP model, which processes each frame individually. This contrasts with UMT, which processes all frame tokens simultaneously. To align with our method, we introduce a new baseline UMT4Clip, where UMT handles frames separately. These differences in token processing strategies inevitably lead to a decline in performance. Following the practice of UMT, we conduct fully fine-tuning experiments on both text-video retrieval and video QA tasks. In Table 7, our TempMe achieves performance close to UMT while requiring only one-third of the GFLOPs and training time. Moreover, we also extend our method to video QA in Table 7. These results demonstrate the superiority and generalizability of our method.
@@
To further validate the generalization capability, we integrate our TempMe with other efficient fine-tuning methods such as Prompt and Adapter. Moreover, we evaluate our TempMe on other backbones. These detailed experimental results are provided in Appendix C.
Table 8: Ablation analysis of individual components. We evaluate in the text-to-video task on MSRVTT with CLIP-ViT-B/32.
@@
CLIP-Straight - 53.0 30.8 53.8 63.3 147.9 41.8
+LoRA 0.49 53.0 43.7 68.9 80.4 193.0 16.0
+ImgMe 0.49 40.2 42.9 68.3 80.2 191.4 16.2
+ClipMe 0.50 34.8 46.1 71.8 80.7 198.6 14.8
Table 9: Ablation analysis of individual components. We evaluate in the text-to-video task on MSRVTT with CLIP-ViT-B/32.
@@
LoRA 0.49 53.0 43.7 68.9 80.4 193.0 16.0
Temporal Modeling 0.50 54.3 46.3 72.1 81.3 199.7 14.8
Token Reduction 0.49 34.7 41.6 67.9 79.3 188.8 16.6
TempMe 0.50 34.8 46.1 71.8 80.7 198.6 14.8
@@ 4.4 Ablation Study
We conduct an in-depth analysis of TempMe from structural and functional perspectives, respectively.
@@
From a functional perspective, TempMe can be categorized into Temporal Modeling and Token Reduction, aimed at improving accuracy and efficiency, respectively. Table 9 demonstrates the impact of each function. (1) The accuracy improvements are attributed to Temporal Modeling, which aggregates clips progressively to enhance spatio-temporal learning. In this framework, the self-attention modules of the early layers is applied to intra-frame tokens in the spatial domain. In the later layers, they operate on tokens across frames in the spatio-temporal domain. (2) The efficiency improvements arise from Token Reduction, which reduces redundancy in the entire framework. In the early layers, it slightly reduces intra-frame tokens to decrease spatial redundancy. In the later layers, it significantly reduces tokens among frames to address large temporal redundancy. (3) Without the support of Temporal Modeling, Token Reduction alone significantly reduces tokens in the spatial domain, leading to a substantial 4.2% decrease in R-sum. However, due to the considerable temporal redundancy, the combination of Temporal Modeling and Token Reduction (TempMe) significantly reduces complexity while achieving high performance.
@@
For a more comprehensive evaluation of our method, we conduct a series of ablation studies, as detailed in Appendix C. First, since the trainable parameters of TempMe using ViT-B/32 include LoRA (0.49M) and clip positional embeddings (0.01M), we perform ablation on the clip positional embeddings of the ClipMe block. Next, we thoroughly conduct an exhaustive analysis of the merging strategy. Following this, we provide a detailed analysis of the intra-clip and cross-clip merging steps within the ClipMe block. Furthermore, given that the number of sampled frames affects the model complexity, we conduct an ablation study on frame sampling.
Figure 4: Qualitative comparisons on MSRVTT with ViT-B/16. Patches that share the same inner and border color are merged together. TempMe merges tokens of similar elements across frames.
@@ 4.5 Qualitative Results
Figure 4 presents visualization results of ToMe and our TempMe. ToMe only merges spatial tokens within individual frames, leaving a large number of redundant tokens across frames, contributing to high computational overhead. In contrast, TempMe merges tokens of similar elements across contiguous frames. For instance, TempMe merges similar body regions of Patrick Star across frames, such as the head, upper body, and lower body. These visualization results show that our TempMe effectively reduces temporal redundancy. More qualitative results are provided in Appendix C.

### v2

@@ 3.2 Temporal Token Merging
To verify the efficiency of our proposed TempMe, we perform a detailed complexity analysis in Table 1. Following common protocols Luo et al. (2022); Yang et al. (2024), the frame length is set to 12 for MSR-VTT and LSMDC, and to 64 for ActivityNet and DiDeMo. The complexity comparison between ToMe and our TempMe for these frame lengths is shown in Table 1. Compared to ToMe, our TempMe reduces tokens by more than 30% for both CLIP-ViT-B/32 and CLIP-ViT-B/16, effectively decreasing model complexity while significantly surpassing accuracy. Particularly, with a 12-frame length and CLIP-ViT-B/16 backbone on MSRVTT, TempMe outputs only 5% of the input tokens, reaches 57% GFLOPs, and achieves a 5.3% R-Sum gain.
@@ 4 Experiments
Table 2: Comparisons in the text-to-video task on MSRVTT. We evaluate GFLOPs only for the video backbone during inference. A detailed comparison of memory usage can be found in Appendix C.7.
@@
Full Fine-tuning CLIP4Clip 123.54 53.0 43.1 70.4 80.8 194.3 16.2
Parameter-Efficient Prompt 0.08 58.2 40.4 66.3 77.3 184.0 16.7
Adapter 0.26 53.1 41.9 69.9 78.7 190.2 14.9
LoRA 0.49 53.0 43.7 68.9 80.4 193.0 16.0
PLEVU 6.35 - 36.7 64.6 76.8 178.1 -
VoPF+C 14.10 58.0 44.6 69.9 80.3 194.8 16.3
DGLL 0.83 67.4 44.7 70.5 79.2 194.4 16.2
DGLT 9.57 >67.4 45.8 69.3 79.4 194.5 16.3
Parameter-Efficient & Inference-Efficient EVIT 0.49 37.2 41.4 69.0 78.1 188.5 16.8
DiffRate 0.49 36.8 41.5 68.6 79.8 189.9 16.3
STA 0.49 35.7 42.6 69.5 78.8 190.9 17.0
ToMe 0.49 40.2 42.9 68.3 80.2 191.4 16.2
TESTA 0.59 40.6 43.7 69.0 79.4 192.1 16.8
@@
Parameter-Efficient MV-Adapter 3.6 >210 46.0 72.0 82.1 200.1 -
RAP 1.06 >210 46.5 73.9 82.0 202.4 12.1
VoP 14.10 246.2 47.7 72.4 82.2 202.3 12.0
DGLL 0.83 251.2 48.3 71.8 80.6 200.7 13.4
@@ 4.1 Experimental Settings
Following common practice, we perform experiments on four widely used benchmarks for text-video retrieval: MSRVTT Xu et al. (2016), ActivityNet Krishna et al. (2017), DiDeMo Anne Hendricks et al. (2017), and LSMDC Rohrbach et al. (2015). Detailed descriptions of these datasets are provided in Appendix B.
@@
[…]
@@
We compare our TempMe with several strong baselines: (1) CLIP4Clip Luo et al. (2022), a classic full fine-tuning method. For a fair comparison, we only consider the parameter-free type that does not introduce additional modules. (2) Prompt Khattak et al. (2023), LoRA Hu et al. (2021), and Adapter Houlsby et al. (2019), which are mainstream parameter-efficient fine-tuning methods. (3) PLEVU Ju et al. (2022), VoP Huang et al. (2023), DGL Yang et al. (2024), MV-Adapter Jin et al. (2024), and RAP Cao et al. (2024), which are methods tailored for parameter-efficient text-video retrieval. (4) EVIT Liang et al. (2022), DiffRate Chen et al. (2023a), and ToMe Bolya et al. (2022), which are token compression methods for image models. (5) STA Ding et al. (2023) and TESTA Ren et al. (2023), which are video token compression methods. For fairness, token compression methods apply LoRA like our TempMe.
@@
Prompt 29.7 216.8 2369 44.3/194.2
Adapter 29.4 211.7 2364 44.9/196.8
LoRA 30.6 211.3 2364 47.3/201.4
VoPF+C 25.0 246.2 2368 47.7/202.3
DGLL 3.3 251.2 2416 48.3/200.7
DiffRate 40.6 138.5 588 47.3/202.4
ToMe 40.8 144.4 924 46.2/200.6
TempMe 45.1 121.4 127 49.0/206.7
@@
CLIP4Clip 211.3 1.00$\times$ 70.1GB 8.69GB 47.3/202.3
+ToMe 144.4 1.35$\times$ 61.6GB 8.55GB 47.2/201.6
+TempMe 121.4 1.57$\times$ 52.7GB 8.55GB 50.9/210.2
@@ 4.2 Comparisons with State-of-the-art Methods
In Table 2, we compare the t2v performance on MSRVTT. Our TempMe achieves significant improvements over previous methods, with a $\textbf{3.8}\%$ R-Sum increase using ViT-B/32 and a $\textbf{4.3}\%$ R-Sum increase using ViT-B/16, while maintaining minimal GFLOPs. Table 3 shows the t2v results on ActivityNet, DiDeMo, and LSMDC, demonstrating that TempMe consistently outperforms state-of-the-art methods. Detailed v2t results are provided in Appendix C.1.
@@ 4.3 Generalization Analysis
Table 6: Application in video foundation models for text-video retrieval. We evaluate full fine-tuning on MSRVTT for the text-to-video task with UMT-B/16-25M.
@@
Table 7: Application in video foundation models for text-video QA. We evaluate full fine-tuning on MSR-QA with UMT-B/16-25M.
@@
We extend our method to video foundation methods, such as UMT Li et al. (2023). TempMe is built upon the text-image CLIP model, which processes each sampled frame individually. This contrasts with UMT, which processes all frame tokens simultaneously. To align with our method, we introduce a new baseline UMT4Clip, where UMT handles frames separately. These differences in token processing strategies inevitably lead to a decline in performance. Following the practice of UMT, we conduct fully fine-tuning experiments on both text-video retrieval and video QA tasks. In Table 7, our TempMe achieves performance close to UMT while requiring only one-third of the GFLOPs and training time. Moreover, we also extend our method to video QA in Table 7. These results demonstrate the generalizability of our method.
@@
To further validate the generalization capability, we integrate our TempMe with other parameter-efficient fine-tuning methods such as Prompt and Adapter in Appendix C.2. Moreover, we evaluate our TempMe on other backbones. These detailed experimental results are provided in Appendix C.3.
Table 8: Ablation on each component in t2v on MSRVTT with CLIP-ViT-B/32. CLIP-straight denotes the zero-shot performance.
@@
CLIP-Straight 53.0 30.8 147.9 41.8
+LoRA 53.0 43.7 193.0 16.0
+ImgMe 40.2 42.9 191.4 16.2
+ClipMe 34.8 46.1 198.6 14.8
Table 9: Ablation on each function in t2v on MSRVTT with CLIP-ViT-B/32. Temporal Modeling is shown in Figure 6 of Appendix.
@@
LoRA 53.0 43.7 193.0 16.0
Temporal Modeling 54.3 46.3 199.7 14.8
Token Reduction 34.7 41.6 188.8 16.6
TempMe 34.8 46.1 198.6 14.8
@@ 4.4 Ablation Study
We conduct an analysis of TempMe from structural and functional perspectives, respectively.
@@
From a functional perspective, TempMe can be categorized into Temporal Modeling and Token Reduction, aimed at improving accuracy and efficiency, respectively. Table 9 demonstrates the impact of each function. (1) The accuracy improvements are attributed to Temporal Modeling, which aggregates clips progressively to enhance spatio-temporal learning. In this framework, the attention modules of the early layers are applied to intra-frame tokens in the spatial domain. In the later layers, they operate on tokens across frames in the spatio-temporal domain. (2) The efficiency improvements arise from Token Reduction, which reduces redundancy in the entire framework. In the early layers, it slightly reduces intra-frame tokens to decrease spatial redundancy. In the later layers, it significantly reduces tokens among frames to address large temporal redundancy. (3) Without the support of Temporal Modeling, Token Reduction alone significantly reduces tokens in the spatial domain, leading to a substantial 4.2% decrease in R-sum. However, due to the considerable temporal redundancy, the combination of Temporal Modeling and Token Reduction (TempMe) significantly reduces complexity while achieving high performance.
Table 10: Ablation analysis of different merging strategies on MSRVTT with CLIP-VIT-B/32. We adopt progressive, holistic, and continuous merging (A0) as our final strategy.
Methods GFLOPs Text-to-Video Video-to-Text
[…]
Ablation on Merging Strategy.
Table 10 validates the effectiveness of the merging strategy utilized in our PMG framework. (1) Progressive Merging. In A1, we observe that merging all clips at once leads to high model complexity, due to the quadratic complexity of MHSA with the surge in token numbers. Alternatively, the progressive merging of clips benefits video-level information. Our A0 with lower complexity achieves better performance. (2) Holistic Merging. We conduct experiments in A2 to evaluate the effectiveness of merging all clips. A2 shows that a notable performance decline when frames are not fully merged (e.g., 12 frames into 3 or 4 frames). These findings indicate that encoding all frame information simultaneously in MHSA is crucial for holistic video understanding. (3) Continuous Merging. Experiments on merging with various gaps are conducted in A3. Larger gaps trigger an earlier start of the clip merging process, which leads to a higher token reduction but also a drop in performance. In conclusion, we adopt progressive, holistic, and continuous merging (A0) as our final strategy.
@@
We further conduct a series of ablation studies. First, since the trainable parameters of TempMe using CLIP-ViT-B/32 include LoRA ($\sim$0.49M) and clip positional embeddings ($\sim$0.01M), we perform ablation on the clip positional embeddings of the ClipMe block in Appendix C.4. Next, we provide a detailed analysis of the intra-clip and cross-clip merging steps within the ClipMe block in Appendix C.5. Furthermore, given that the number of sampled frames affects the model complexity, we conduct an ablation study on frame sampling in Appendix C.6. Finally, we report memory usage during training and inference in Appendix C.7.
Figure 4: Qualitative comparisons on MSRVTT with CLIP-ViT-B/16. Patches that share the same inner and border color are merged. TempMe merges tokens of similar elements across frames.
@@ 4.5 Qualitative Results
Figure 4 presents visualization results of ToMe and our TempMe. ToMe only merges spatial tokens within individual frames, leaving a large number of redundant tokens across frames, contributing to high computational overhead. In contrast, TempMe merges tokens of similar elements across contiguous frames. For instance, TempMe merges similar body regions of Patrick Star across frames, such as the head, upper body, and lower body. These visualization results show that our TempMe effectively reduces temporal redundancy. More qualitative results are provided in Appendix C.8.